# Logarithmic critical slowing down in complex systems: from statics to dynamics

### v3

@@ I Introduction
Hereafter, we present a general method for the quantitative computation of the coefficient of the logarithmic decay of the density-density correlation functions in viscous liquids. Göetze and Sjögren Gotze and Sjogren (1989) predicted a $1/\ln t^{2}$ decay exactly at the $A_{3}$ singularity and a behavior of the kind $-(\ln t)^{\gamma}$ as $a\sim 0$, $\lambda\sim 1$. Their exemplifying case is the $F_{13}$ mode-coupling schematic theory Götze (2009). However, in this case, the $A_{3}$ singularity cannot be directly accessed in experiments or in numerical simulations because it occurs in the region of the phase diagram pertaining to the glassy phase, beyond the dynamic arrest line where TTI breaks down and the MCT does not hold anymore. Therefore, in the liquid phase the presence of the $A_{3}$ singularity is only felt in weakly logarithmic corrections to the power-law $\beta$ decay in regions of the space parameters (temperature, packing fraction, …) close enough to it. Other systems displaying this kind of singularity include disordered spin-glass models Crisanti et al. (1993a); Crisanti and Leuzzi (2006, 2015); Franz et al. (2013); Cammarota and Biroli (2012a), liquids in porous media both in the MCT Krakoviack (2005, 2007) and in the hypernetted-chain approximations Madden and Glandt (1988); Given and Stell (1992) and liquid models with pinned particles Cammarota and Biroli (2012b). The $-\ln t$ behavior of the correlation function appears to be the correct fitting law for about a decade or two in most of the known experiments and numerical simulations of repulsive colloids Sjogren (1991); Zaccarelli et al. (2002); Sciortino et al. (2003a, b); Götze (2009). Also cases where the $A_{3}$ singularity is directly accessible from the liquid phase are devised in MCT, for instance in the $F_{12}$ schematic model Götze (2009). The quantitative estimation of the parameters of the logarithmic behavior can, so far, be performed exactly only in MCT schematic theories. Moreover, as, e.g., in the case of the $F_{13}$ model, the $A_{3}$ singularities lie in the region where one of the fundamental assumptions on which MCT is built, TTI, does not hold.
@@
In many systems the replica method offers a way to characterize dynamical arrest phenomena in a purely static framework which is often simpler than a dynamical approach. It is, therefore, natural to assume that universal static critical properties can be obtained a la Landau from simple assumptions on the (replicated) Gibbs free energy at the corresponding critical point. A decade ago, it was realized that replicated theories also determine important features of critical glassy dynamics Caltagirone et al. (2012a, b, c); Ferrari et al. (2012); Parisi and Rizzo (2013a). Notably, they give the same scale-invariant equations for the critical correlators that are often obtained by studying the actual dynamical equations of these systems. An important consequence is that the exponent parameter $\lambda$, Eq. (1), can be computed in a static replicated theory.
@@ II Outline of the Results
We will follow and expand the derivation of Ref. Parisi and Rizzo (2013a), considering theories in which the argument of the Gibbs free energy $G(Q)$ is a replicated matrix $Q_{ab}$ with $a=1,\dots,n$. The replica number $n$ in these theories is usually continued from integer to real continuous values, and we will take into account the two important cases $n\rightarrow 0$ and $n\rightarrow 1$. Such $Q_{ab}$ matrix naturally appears in Spin-Glasses where, due to the presence of quenched disorder, one resorts for technical reasons to the replica method. From a physical point of view, the order parameter to identify a glassy phase, i.e., a ”multi-equilibria” phase composed by many different states, cannot rely on an absolute reference for a state, since no a priori clear pattern is provided because of frustration. As a consequence, the order parameter is built on the similarity between different states. More precisely, on the whole range (hierarchy) of possible similarities, summed up in a probability distribution for the values of the overlap matrix elements. In this case $Q_{ab}$ is naturally identified with the average of the overlap between two different replicas of the system
@@
The above expression can be obtained from a microscopic description in a variety of contexts Mézard et al. (1987); Parisi et al. (2020). In the above expression we have retained only the terms relevant for the present discussion (the $y_{3}$ term, actually, vanishes, as will be shown in Sec. III.1), while the complete expression has actually eight third-order terms and twenty-three fourth-order terms that will be displayed later, in appendix A. At the end of section IV we will explain why the other terms can be neglected. We will focus on critical points characterized by the condition $m_{1}=0$. Depending on the values of the remaining parameters and on the replica number $n$ we may have different types of transition. Three such transitions, discussed in detail in Parisi and Rizzo (2013a) are:
@@
    $m_{2}\neq 0\neq m_{3}$ and replica number $n\to 0$, that corresponds to the SG transition in a field that occurs along the de Almeida-Thouless line Mézard et al. (1987); Crisanti et al. (1993b), and
@@
In the liquid/paramagnetic phase the function $C(t)$ decays exponentially but the correlation time diverges at the critical point. As mentioned in the introduction it has been shown Parisi and Rizzo (2013a) that the structure of the replicated Gibbs free energy at the critical point determines also the essential features of critical dynamics. More precisely in the case of the SG transition (i) one can show that the TTI correlation at large time-differences $t$ is described by:
@@
where the $J^{\prime}s$ are quenched random interactions and the $s_{i}$ can be either Ising spin or satisfy a spherical constraint. In the spherical $3$-spin case in presence of a magnetic field there is a tricritical point with $\lambda=1$ in the temperature-magnetic field plane where a line of discontinuous transitions meets a line of continuous transitions Crisanti and Sommers (1992); Crisanti et al. (1993a). Another example is that of a mixed $2+3$ model that corresponds to the so-called schematic $F_{12}$ model in the context of MCT. In the phase diagram, e.g., in the plane of the magnitudes of the $2$-spin and the $3$-spin interaction there is a $\lambda=1$ critical point. Upon increasing the relative magnitude of the $3$-spin interaction, a line of continuous transitions meets a line of discontinuous transitions Franz et al. (2013); Caltagirone et al. (2013). Random pinning of a spherical $p$-spin glass model, i.e. freezing a fraction $c$ of the spins Cammarota and Biroli (2012b, a), is also relevant: in the temperature-concentration plane there is a line of discontinuous transitions that, upon increasing the concentration, ends in a point characterized by $\lambda=1$. Finally we mention the Potts Spin-Glass with Hamiltonian:
@@ III Derivation of the Equations of Critical Dynamics
Now we translate the above equation into an equation for the dynamical correlation valid at large times and near the critical point. We will briefly sketch the arguments leading to this mapping but we refer the reader to Ref. Parisi and Rizzo (2013a) for all the details of the procedure. The result is obtained in the context of a super-field formulation of dynamics Kurchan (1992) in which both the dynamical correlation and response functions are represented by a single dynamical order parameter $Q(1,2)$ in terms of (commuting) times $t_{1,2}$ and Grassmannian anticommuting variables $\theta_{1,2}$, $\bar{\theta}_{1,2}$.
@@ III.1 Theories with $n=0$
We note that this representation is appropriate for the phase transitions characterized by a replicated free energy with $n=0$, whereas for $n=1$ a different representation must be considered (Parisi and Rizzo (2013a), Sec. III.D). We postpone the discussion of this case to the end of this section. On general grounds it is to be expected that the dynamical order parameter $Q(1,2)$ at large times must be related to the static order parameter $Q_{ab}$. Indeed, as noted in Kurchan (1992), the static result is obtained in the so-called Fast Motion (FM) limit that corresponds to an infinitely fast microscopic dynamics. In this limit, configurations at different times are completely uncorrelated and they are equivalent to different replicas of the same system. As a consequence, in this limit $Q_{FM}(1,2)$ has a diagonal structure:
@@
The dynamical equations for $Q(1,2)$ can be obtained from a dynamical Gibbs free energy and one may expect that the critical dynamics is determined a la Landau by its expansion in powers of $\delta Q(1,2)$. In Parisi and Rizzo (2013a) it is argued that the dynamical Gibbs free energy must have the same structure of the replicated Gibbs free energy with the same coupling constants and, therefore, Eq. (III) translates into an identical equation for $\delta Q(1,2)$.
@@ III.2 Theories with $n=1$
We now turn to the $n=1$ case. This is relevant for the important case of the dynamic transition in SG models with one step of Replica-Symmetry-Breaking and in glass-forming supercooled liquids Crisanti (2008); Franz et al. (2011); Parisi and Rizzo (2013a); Rizzo (2016); Parisi et al. (2020). As we mentioned before, here the dynamics has to be described in a formalism that takes into account the initial condition. The formulation is given in section III.D of Parisi and Rizzo (2013a) and a complete treatment is reported in Rizzo (2016). We will not enter into the details hereafter and we will limit ourselves to mention that the dynamical objects involved are, once again, the super-symmetric matrices $Q(1,2)$. Inasmuch as we did above, these can be written in terms of a single TTI correlation function $C(t)=C(-t)$:
@@
Once again Eq. (III) can be, thus, translated into an equation for $C(t)$. As we saw before we only need to specify the analogue for $n=1$ of Eqs. (34, 35), i.e., the behavior of element-wise products $A^{k}(1,2)$ and dot products $\int A(1,3)B(3,2)\,d3$. For the products we just have Parisi and Rizzo (2013a)
@@ IV Expansion of the Replicated Free Energy at Fourth Order
In this section we express the coefficients of the free energy of a generic model in terms of physical observables, namely the cumulants of the overlap. For the sake of clarity in the following we will reproduce the derivation of the third order expansion, initially reported in Ref. Parisi and Rizzo (2013b), while we will postpone to appendix A the derivation of the fourth order result. For the sake of readability we might repeat some of the definitions already given before.
@@
We note that the above free energy functional arises if we apply to each spin $s_{i}^{a}$ of each replica a Gaussian distributed random field $h_{i}^{a}$ with covariance matrix given by $\overline{h_{i}^{a}h_{j}^{b}}=\lambda_{ab}\delta_{ij}$. Following Parisi and Rizzo (2013b) we start by expanding $F(\lambda)$ in powers of $\lambda$ at fourth order assuming $\lambda_{aa}=0$ afor-$\forall a$:
@@ VI Conclusions
It should be also noted that we have been considering equilibrium dynamics, therefore the results are limited to the temperature regime in which the system can still be thermalized and there is no aging. Nonetheless it is not unlikely that, with more effort, the analysis can be extended to the aging regime, as it is known that static replicated theories can be connected to off-equilibrium dynamics as well Cugliandolo and Kurchan (1993); Caltagirone et al. (2012d).

### v4

@@ I Introduction
Hereafter, we present a general method for the quantitative computation of the coefficient of the logarithmic decay of the density-density correlation functions in viscous liquids. Göetze and Sjögren Gotze and Sjogren (1989) predicted a $1/\ln t^{2}$ decay exactly at the $A_{3}$ singularity and a behavior of the kind $-(\ln t)^{\gamma}$ as $a\sim 0$, $\lambda\sim 1$. Their exemplifying case is the $F_{13}$ mode-coupling schematic theory Götze (2009). However, in this case, the $A_{3}$ singularity cannot be directly accessed in experiments or in numerical simulations because it occurs in the region of the phase diagram pertaining to the glassy phase, beyond the dynamic arrest line where TTI breaks down and the MCT does not hold anymore. Therefore, in the liquid phase the presence of the $A_{3}$ singularity is only felt in weakly logarithmic corrections to the power-law $\beta$ decay in regions of the space parameters (temperature, packing fraction, …) close enough to it. Other systems displaying this kind of singularity include disordered spin-glass models Crisanti et al. (1993); Crisanti and Leuzzi (2006, 2015); Franz et al. (2013); Cammarota and Biroli (2012a), liquids in porous media both in the MCT Krakoviack (2005, 2007) and in the hypernetted-chain approximations Madden and Glandt (1988); Given and Stell (1992) and liquid models with pinned particles Cammarota and Biroli (2012b). The $-\ln t$ behavior of the correlation function appears to be the correct fitting law for about a decade or two in most of the known experiments and numerical simulations of repulsive colloids Sjogren (1991); Zaccarelli et al. (2002); Sciortino et al. (2003a, b); Götze (2009). Also cases where the $A_{3}$ singularity is directly accessible from the liquid phase are devised in MCT, for instance in the $F_{12}$ schematic model Götze (2009). The quantitative estimation of the parameters of the logarithmic behavior can, so far, be performed exactly only in MCT schematic theories. Moreover, as, e.g., in the case of the $F_{13}$ model, the $A_{3}$ singularities lie in the region where one of the fundamental assumptions on which MCT is built, TTI, does not hold.
@@
In many systems the replica method offers a way to characterize dynamical arrest phenomena in a purely static framework which is often simpler than a dynamical approach. It is, therefore, natural to assume that universal static critical properties can be obtained a la Landau from simple assumptions on the (replicated) Gibbs free energy at the corresponding critical point. A decade ago, it was realized that replicated theories also determine important features of critical glassy dynamics Caltagirone et al. (2012a, b, c); Ferrari et al. (2012); Parisi and Rizzo (2013). Notably, they give the same scale-invariant equations for the critical correlators that are often obtained by studying the actual dynamical equations of these systems. An important consequence is that the exponent parameter $\lambda$, Eq. (1), can be computed in a static replicated theory.
@@ II Outline of the Results
We will follow and expand the derivation of Ref. Parisi and Rizzo (2013), considering theories in which the argument of the Gibbs free energy $G(Q)$ is a replicated matrix $Q_{ab}$ with $a=1,\dots,n$. The replica number $n$ in these theories is usually continued from integer to real continuous values, and we will take into account the two important cases $n\rightarrow 0$ and $n\rightarrow 1$. Such $Q_{ab}$ matrix naturally appears in Spin-Glasses where, due to the presence of quenched disorder, one resorts for technical reasons to the replica method. From a physical point of view, the order parameter to identify a glassy phase, i.e., a ”multi-equilibria” phase composed by many different states, cannot rely on an absolute reference for a state, since no a priori clear pattern is provided because of frustration. As a consequence, the order parameter is built on the similarity between different states. More precisely, on the whole range (hierarchy) of possible similarities, summed up in a probability distribution for the values of the overlap matrix elements. In this case $Q_{ab}$ is naturally identified with the average of the overlap between two different replicas of the system
@@
The above expression can be obtained from a microscopic description in a variety of contexts Mézard et al. (1987); Parisi et al. (2020). In the above expression we have retained only the terms relevant for the present discussion (the $y_{3}$ term, actually, vanishes, as will be shown in Sec. III.1), while the complete expression has actually eight third-order terms and twenty-three fourth-order terms that will be displayed later, in appendix A. At the end of section IV we will explain why the other terms can be neglected. We will focus on critical points characterized by the condition $m_{1}=0$. Depending on the values of the remaining parameters and on the replica number $n$ we may have different types of transition. Three such transitions, discussed in detail in Parisi and Rizzo (2013) are:
@@
    $m_{2}\neq 0\neq m_{3}$ and replica number $n\to 0$, that corresponds to the SG transition in a field that occurs along the de Almeida-Thouless line Mézard et al. (1987); Crisanti et al. (1993), and
@@
In the liquid/paramagnetic phase the function $C(t)$ decays exponentially but the correlation time diverges at the critical point. As mentioned in the introduction it has been shown Parisi and Rizzo (2013) that the structure of the replicated Gibbs free energy at the critical point determines also the essential features of critical dynamics. More precisely in the case of the SG transition (i) one can show that the TTI correlation at large time-differences $t$ is described by:
@@
where the $J^{\prime}s$ are quenched random interactions and the $s_{i}$ can be either Ising spin or satisfy a spherical constraint. In the spherical $3$-spin case in presence of a magnetic field there is a tricritical point with $\lambda=1$ in the temperature-magnetic field plane where a line of discontinuous transitions meets a line of continuous transitions Crisanti and Sommers (1992); Crisanti et al. (1993). Another example is that of a mixed $2+3$ model that corresponds to the so-called schematic $F_{12}$ model in the context of MCT. In the phase diagram, e.g., in the plane of the magnitudes of the $2$-spin and the $3$-spin interaction there is a $\lambda=1$ critical point. Upon increasing the relative magnitude of the $3$-spin interaction, a line of continuous transitions meets a line of discontinuous transitions Franz et al. (2013); Caltagirone et al. (2013). Random pinning of a spherical $p$-spin glass model, i.e. freezing a fraction $c$ of the spins Cammarota and Biroli (2012b, a), is also relevant: in the temperature-concentration plane there is a line of discontinuous transitions that, upon increasing the concentration, ends in a point characterized by $\lambda=1$. Finally we mention the Potts Spin-Glass with Hamiltonian:
@@ III Derivation of the Equations of Critical Dynamics
Now we translate the above equation into an equation for the dynamical correlation valid at large times and near the critical point. We will briefly sketch the arguments leading to this mapping but we refer the reader to Ref. Parisi and Rizzo (2013) for all the details of the procedure. The result is obtained in the context of a super-field formulation of dynamics Kurchan (1992) in which both the dynamical correlation and response functions are represented by a single dynamical order parameter $Q(1,2)$ in terms of (commuting) times $t_{1,2}$ and Grassmannian anticommuting variables $\theta_{1,2}$, $\bar{\theta}_{1,2}$.
@@ III.1 Theories with $n=0$
We note that this representation is appropriate for the phase transitions characterized by a replicated free energy with $n=0$, whereas for $n=1$ a different representation must be considered (Parisi and Rizzo (2013), Sec. III.D). We postpone the discussion of this case to the end of this section. On general grounds it is to be expected that the dynamical order parameter $Q(1,2)$ at large times must be related to the static order parameter $Q_{ab}$. Indeed, as noted in Kurchan (1992), the static result is obtained in the so-called Fast Motion (FM) limit that corresponds to an infinitely fast microscopic dynamics. In this limit, configurations at different times are completely uncorrelated and they are equivalent to different replicas of the same system. As a consequence, in this limit $Q_{FM}(1,2)$ has a diagonal structure:
@@
The dynamical equations for $Q(1,2)$ can be obtained from a dynamical Gibbs free energy and one may expect that the critical dynamics is determined a la Landau by its expansion in powers of $\delta Q(1,2)$. In Parisi and Rizzo (2013) it is argued that the dynamical Gibbs free energy must have the same structure of the replicated Gibbs free energy with the same coupling constants and, therefore, Eq. (III) translates into an identical equation for $\delta Q(1,2)$.
@@ III.2 Theories with $n=1$
We now turn to the $n=1$ case. This is relevant for the important case of the dynamic transition in SG models with one step of Replica-Symmetry-Breaking and in glass-forming supercooled liquids Crisanti (2008); Franz et al. (2011); Parisi and Rizzo (2013); Rizzo (2016); Parisi et al. (2020). As we mentioned before, here the dynamics has to be described in a formalism that takes into account the initial condition. The formulation is given in section III.D of Parisi and Rizzo (2013) and a complete treatment is reported in Rizzo (2016). We will not enter into the details hereafter and we will limit ourselves to mention that the dynamical objects involved are, once again, the super-symmetric matrices $Q(1,2)$. Inasmuch as we did above, these can be written in terms of a single TTI correlation function $C(t)=C(-t)$:
@@
Once again Eq. (III) can be, thus, translated into an equation for $C(t)$. As we saw before we only need to specify the analogue for $n=1$ of Eqs. (34, 35), i.e., the behavior of element-wise products $A^{k}(1,2)$ and dot products $\int A(1,3)B(3,2)\,d3$. For the products we just have Parisi and Rizzo (2013)
@@ IV Expansion of the Replicated Free Energy at Fourth Order
In this section we express the coefficients of the free energy of a generic model in terms of physical observables, namely the cumulants of the overlap. For the sake of clarity in the following we will reproduce the derivation of the third order expansion, initially reported in Ref. Parisi and Rizzo (2013), while we will postpone to appendix A the derivation of the fourth order result. For the sake of readability we might repeat some of the definitions already given before.
@@
We note that the above free energy functional arises if we apply to each spin $s_{i}^{a}$ of each replica a Gaussian distributed random field $h_{i}^{a}$ with covariance matrix given by $\overline{h_{i}^{a}h_{j}^{b}}=\lambda_{ab}\delta_{ij}$. Following Parisi and Rizzo (2013) we start by expanding $F(\lambda)$ in powers of $\lambda$ at fourth order assuming $\lambda_{aa}=0$ afor-$\forall a$:
@@ VI Conclusions
It should be also noted that we have been considering equilibrium dynamics, therefore the results are limited to the temperature regime in which the system can still be thermalized and there is no aging. Nonetheless it is not unlikely that, with more effort, the analysis can be extended to the aging regime, as it is known that static replicated theories can be connected to off-equilibrium dynamics as well Cugliandolo and Kurchan (1993); caltagirone12c.